OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (anonymous):

y-4=2/3x-2 y=2/3x+2 Because it's parallel, it will have the same slope. y-y1=m(x-x1) This is standard form. y-(-2)=2/3(x-1) y+2=2/3(x-1) or y=2/3x-8/3

OpenStudy (anonymous):

cant it be 3y+6=2x-2 ?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

wops i mean -2x+3y=-8